# AED: Adaptable Error Detection for Few-shot Imitation Policy

### v1

@@ Abstract
We study how to report few-shot imitation (FSI) policies’ behavior errors in novel environments, a novel task named adaptable error detection (AED). The potential to cause serious damage to surrounding areas limits the application of FSI policies in real-world scenarios. Thus, a robust system is necessary to notify operators when FSI policies are inconsistent with the intent of demonstrations. We develop a cross-domain benchmark for the challenging AED task, consisting of 329 base and 158 novel environments. This task introduces three challenges, including (1) detecting behavior errors in novel environments, (2) behavior errors occurring without revealing notable changes, and (3) lacking complete temporal information of the rollout due to the necessity of online detection. To address these challenges, we propose Pattern Observer (PrObe) to parse discernible patterns in the policy feature representations of normal or error states, whose effectiveness is verified in the proposed benchmark. Through our comprehensive evaluation, PrObe consistently surpasses strong baselines and demonstrates a robust capability to identify errors arising from a wide range of FSI policies. Moreover, we conduct comprehensive ablations and experiments (error correction, demonstration quality, etc.) to validate the practicality of our proposed task and methodology.
Few-shot Imitation, Adaptable Error Detection, AED
@@ 1 Introduction
Few-shot imitation (FSI), a framework that learns a policy in novel (unseen) environments from a few demonstrations, has recently drawn significant attention in the community (Dance et al., 2021; Hakhamaneshi et al., 2022; Tao et al., 2023; Shin et al., 2023; Liu et al., 2023b; Jiang et al., 2023). Notably, the framework, as exemplified by (Duan et al., 2017; James et al., 2018; Bonardi et al., 2020; Dasari & Gupta, 2020; Dance et al., 2021; Yeh et al., 2022; Hakhamaneshi et al., 2022), has demonstrated its efficacy across a range of robotic manipulation tasks. This framework shows significant potential to adapt to a new task based on just a few demonstrations from their owners (Dance et al., 2021; Hakhamaneshi et al., 2022). However, a major barrier that still limits their ability to infiltrate our everyday lives is the ability to detect behavior errors in novel environments.
We propose a challenging and crucial task called adaptable error detection (AED), aiming to monitor FSI policies and report their behavior errors, and the corresponding benchmark. In this work, behavior errors refer to states that deviate from the demonstrated behavior, necessitating the timely termination of the policy upon their occurrence. Unlike existing few-shot visual perception tasks (Song et al., 2023), the failure can result in significant disruptions to surrounding objects and humans in the real world. Due to this nature, this often restricts real-world experiments to simple tasks. The AED benchmark we propose is built within Coppeliasim (Robotics, ) and Pyrep (James et al., 2019), encompassing six indoor tasks and one factory task. This comprehensive benchmark comprises 329 base environments and 158 new environments, spanning diverse domains and incorporating multiple stages. We aim to create a thorough evaluation platform for AED methodologies, ensuring their effectiveness before real-world deployment.
@@
The AED tasks present three novel challenges, illustrated in Figure 1. First, AED entails monitoring a policy’s behavior in novel environments, the normal states of which are not observed during training. Second, detecting behavior errors becomes challenging, as there are no noticeable changes to indicate when such errors occur. Specifically, this makes it difficult to discern either minor visual differences or misunderstandings of the task through adjacent frames. Third, AED requires online detection to terminate the policy timely, lacking complete temporal information of the rollout.
@@
Current approaches struggle to address the unique challenges posed by AED. One-class classification (OCC) methods, including one-class SVMs (Zhou & Paffenroth, 2017) or autoencoders (Ruff et al., 2018; Park et al., 2018; Chen et al., 2020; Wong et al., 2021), face difficulties in handling unseen environments. These methods are trained solely with normal samples and identify anomalies by evaluating their significant deviation from the learned distribution. However, normal states in novel environments, which include unseen backgrounds and objects, are already considered out-of-distribution for these techniques, resulting in subpar performance. Multiple instance learning (Maron & Lozano-Pérez, 1997) or patch-based methods (Sultani et al., 2018; Roth et al., 2022) may alleviate the second challenge, particularly minor visual differences, in seen environments. However, the feasibility of applying these methods to AED remains underexplored. Video few-shot anomaly detection (vFSAD) methods (Sultani et al., 2018; Feng et al., 2021; Doshi & Yilmaz, 2023) are unsuitable for AED due to their dependency on full temporal observation from videos.
@@
To this end, we introduce Pattern Observer (PrObe), a novel algorithm that extracts discriminative features from the monitored policy to identify instances of behavior errors. Specifically, PrObe designs a gating mechanism to extract task-relevant features from the monitored FSI policy to mitigate the impact of a novel environment. Then, we design an effective loss function to distill sparse pattern features, making it easier to observe changes in observation. Additionally, we design a recurrent generator to generate a pattern flow of current policy. Finally, we determine if there is a behavior error by proposing a novel temporal-aware contrastive objective to compare the pattern flow and demonstrations.
@@
We make three significant contributions in this study: (1) We define a vital yet underexplored task called adaptable error detection (AED). (2) We introduce a novel benchmark and present PrObe, facilitating the research community to delve into this important area collectively. (3) We conduct thorough evaluations on our proposed benchmark, demonstrating the effectiveness of PrObe. It surpasses several baseline models and exhibits robustness across different policies. Our research is anticipated to serve as a key foundation for future real-world experiments in the field of FSI research.
@@ 2.1 Few-shot Imitation (FSI)
With the recent progress in meta-learning (Finn et al., 2017a), the community explores the paradigm for learning policies from limited demonstrations during inference (Finn et al., 2017b; Yu et al., 2018, 2019a). Notably, these works either assume that the agent and expert have the same configuration or explicitly learn a motion mapping between them (Yu et al., 2018). Conversely, DC methods (Dance et al., 2021; Yeh et al., 2022; Watahiki & Tsuruoka, 2022) develop a policy that behaves conditioned both on the current state and demonstrations. Furthermore, they implicitly learn the mapping between agents and experts, making fine-tuning optional. Thereby, effectively extracting knowledge from demonstrations becomes the most critical matter. In most FSI works, no environment interactions are allowed before inference. Hence, policies are usually trained by behavior cloning (BC) objectives, i.e., learning the likelihood of expert actions by giving expert observations. Recently, DCRL (Dance et al., 2021) trains a model using reinforcement learning (RL) objects and performs FSI tasks without interacting with novel environments.
@@
Since humans can perform complex long-horizon tasks after watching a few demonstrations, FSI studies continuously pursue solving long-horizon tasks to verify if machines can achieve the same level. A set of research applies policy on a robot arm to perform daily life tasks in the real world (Yu et al., 2018) or simulation (Duan et al., 2017). The task is usually multi-stage and composed of primitives/skills/stages (Yeh et al., 2022; Hakhamaneshi et al., 2022), such as a typical pick-and-place task (Yu et al., 2019a) or multiple boxes stacking (Duan et al., 2017). Besides, MoMaRT (Wong et al., 2021) tackles a challenging mobile robot task in a kitchen scene.
@@ 2.2 Few-shot Anomaly Detection (FSAD)
Most existing FSAD studies deal with anomalies in images (Roth et al., 2022; Sheynin et al., 2021; Pang et al., 2021; Huang et al., 2022a; Wang et al., 2022) and only a few tackle anomalies in videos (Lu et al., 2020; Qiang et al., 2021; Huang et al., 2022b; Doshi & Yilmaz, 2023). Moreover, problem settings are diverse. Some works presume only normal data are given during training (Sheynin et al., 2021; Wang et al., 2022; Lu et al., 2020; Qiang et al., 2021), while others train models with normal and a few anomaly data and include unknown anomaly classes during inference (Pang et al., 2021; Huang et al., 2022b).
@@
Studies that only use normal training samples usually develop a reconstruction-based model with auxiliary objectives, such as meta-learning (Lu et al., 2020), optical flow (Qiang et al., 2021), and adversarial training (Sheynin et al., 2021; Qiang et al., 2021). Besides, patch-based methods (Pang et al., 2021; Roth et al., 2022) reveal the performance superiority on main image FSAD benchmark (Bergmann et al., 2019) since the anomaly are tiny defeats. Regarding video few-shot anomaly detection (vFSAD), existing works access a complete video to compute the temporal information for determining if it contains anomalies. In addition, vFSAD benchmarks (Mahadevan et al., 2010; Lu et al., 2013) provide frame-level labels to evaluate the accuracy of locating anomalies in videos.
@@
Although both the vFSAD and our AED task require methods to perform in unseen environments, there are differences: (1) The anomalies in vFSAD involve common objects, while AED methods monitor the policy’s behavior errors. (2) An anomaly in vFSAD results in a notable change in the perception field, such as a cyclist suddenly appearing on the sidewalk (Mahadevan et al., 2010). However, no notable change is evident when a behavior error occurs in AED. (3) The whole video is accessible in vFSAD (Sultani et al., 2018; Feng et al., 2021; Doshi & Yilmaz, 2023), allowing for the leverage of its statistical information. In contrast, AED requires online detection to terminate the policy timely, lacking the complete temporal information of the rollout. These differences make our AED task more challenging and also render vFSAD methods infeasible.
@@ 3 Preliminaries
FSI is a framework worthy of attention, accelerating the development of human-robot collaboration, such as robots completing pre-tasks alone or interacting with humans for highly complex tasks. Following Yeh et al. (2022), a FSI task is associated with a set of base environments $E^{b}$ and novel environments $E^{n}$. In each novel environment $e^{n}\in E^{n}$, a few demonstrations $\mathcal{D}^{n}$ are given. The objective is to seek a policy $\pi$ that achieves the best performance (e.g., success rate) in novel environments leveraging a few demonstrations. Notably, the task in base and novel environments are semantically similar, but their backgrounds and interactive objects are disjoint. Besides, observations and demonstrations are RGB-D images that only provide partial information, as seen in recent studies (Dasari & Gupta, 2020; Yeh et al., 2022). Besides, addressing FSI tasks typically involves three challenges in practice: (1) The task is multi-stage, (2) Demonstrations are length-variant and misaligned, and (3) The expert and agent have a distinct appearance or configuration, as presented in Yeh et al. (2022). Developing a policy to solve the FSI task and simultaneously tackle these challenges is crucial.
@@
Figure 2: AED protocol. Before training, we collect successful agent rollouts $X^{b}_{succ}$, failed agent rollouts $X^{b}_{fail}$, and a few expert demonstrations $\mathcal{D}^{b}$ for all base environments $E^{b}$. Then, the process contains three phases: policy training, AED training, and AED inference. The adaptable error detector $\phi$ aims to report policy $\pi_{\theta}$’s behavior errors when performing in novel environments $E^{n}$.
@@
4 AED task
Our AED task is formulated to monitor and report FSI policies’ behavior errors, i.e., states of policy rollouts that are inconsistent with demonstrations. The challenges associated with this task have been presented in Section 1. Next, we formally state our novel AED task.
Problem statement
Let $c$ denote the category of agent’s behavior status, where $c\in\mathbf{C},\mathbf{C}=\{\textrm{normal},\textrm{error}\}$. When the agent with the trained FSI policy $\pi_{\theta}$ performs in a novel environment $e^{n}$, an adaptable error detector $\phi$ can access the agent’s rollout history $h$ and few expert demonstrations $\mathcal{D}^{n}$. It then predicts the categorical probability $\hat{y}$ of the behavior status for the latest state in the history $h$ by
$\hat{y}=\phi(h,\mathcal{D}^{n})=P(c\mid enc(h,\mathcal{D}^{n})),$ (1)
where $enc$ denotes the feature encoder, and it may be $enc_{\phi}$ ($\phi$ contains its own encoder) or $enc_{\pi_{\theta}}$ (based on policy’s encoder). Next, let $y$ represent the ground truth probability, we evaluate $\phi$ via the expectation of detection accuracy over agent rollouts $X^{n}$ in all novel environments $E^{n}$:
@@ 4 AED task
where $A(\cdot,\cdot)$ is the accuracy function that returns 1 if $\hat{y}$ is consistent with $y$ and 0 otherwise. However, frame-level labels are lacking in novel environments since we only know if the rollout is successful or failed at the end. Hence, we present the details of $A(\cdot,\cdot)$ in the Appendix’s Section E.1.
@@
To train the detector $\phi$, we assume it can access the resources in base environments $E^{b}$ just like FSI policies, i.e., successful agent rollouts $X^{b}_{succ}$ and a few expert demonstrations $\mathcal{D}^{b}$, which are collected by executing the agent’s and expert’s optimal policies $\pi^{\ast}$ [^1] The agent’s optimal policy is expected to be known during training. Besides, successful agent rollouts are not expert demonstrations since the agent and expert have different configurations., respectively. Besides, we collect failed agent rollouts $X^{b}_{fail}$ by intervening in the agent’s $\pi^{\ast}$ at a critical timing (e.g., the moment to grasp objects) so that $X^{b}_{fail}$ can possess precise frame-level error labels. Figure 2 shows the protocol: (1) Policy $\pi_{\theta}$ is optimized using successful agent rollouts $X^{b}_{succ}$ and few demonstrations $\mathcal{D}^{b}$. (2) Detector $\phi$ is trained on agent rollouts $X^{b}_{succ}$, $X^{b}_{fail}$ and few demonstrations $\mathcal{D}^{b}$. Besides, $\phi$ extracts features from policy $\pi_{\theta}$ if it depends on policy’s encoder. And $\pi_{\theta}$’s parameters are not updated in this duration. (3) No agent rollouts are pre-collected in novel environments $E^{n}$; policy $\pi_{\theta}$ solves the task leveraging few demonstrations $\mathcal{D}^{n}$. And the detector $\phi$ monitors policy $\pi_{\theta}$’s behavior during the progress. Our AED task aims to evaluate the detector $\phi$’s ability to report policy $\pi_{\theta}$’s behavior errors.

### v2

@@ Abstract
We introduce a new task called Adaptable Error Detection (AED), which aims to identify behavior errors in few-shot imitation (FSI) policies based on visual observations in novel environments. The potential to cause serious damage to surrounding areas limits the application of FSI policies in real-world scenarios. Thus, a robust system is necessary to notify operators when FSI policies are inconsistent with the intent of demonstrations. This task introduces three challenges: (1) detecting behavior errors in novel environments, (2) identifying behavior errors that occur without revealing notable changes, and (3) lacking complete temporal information of the rollout due to the necessity of online detection. However, the existing benchmarks cannot support the development of AED because their tasks do not present all these challenges. To this end, we develop a cross-domain AED benchmark, consisting of 322 base and 153 novel environments. Additionally, we propose Pattern Observer (PrObe) to address these challenges. PrObe is equipped with a powerful pattern extractor and guided by novel learning objectives to parse discernible patterns in the policy feature representations of normal or error states. Through our comprehensive evaluation, PrObe demonstrates superior capability to detect errors arising from a wide range of FSI policies, consistently surpassing strong baselines. Moreover, we conduct detailed ablations and a pilot study on error correction to validate the effectiveness of the proposed architecture design and the practicality of the AED task, respectively.
@@ 1 Introduction
Few-shot imitation (FSI), a framework that learns a policy in novel (unseen) environments from a few demonstrations, has recently drawn significant attention in the community [1, 2, 3, 4, 5, 6]. Notably, the framework, as exemplified by [7, 8, 9, 10, 1, 11, 2], has demonstrated its efficacy across a range of robotic manipulation tasks. This framework shows significant potential to adapt to a new task based on just a few demonstrations from their owners [1, 2]. However, a major barrier that still limits their ability to infiltrate our everyday lives is the ability to detect behavior errors in novel environments.
We propose a challenging and crucial task called adaptable error detection (AED), aiming to monitor FSI policies from visual observations and report their behavior errors, along with the corresponding benchmark. In this work, behavior errors refer to states that deviate from the demonstrated behavior, necessitating the timely termination of the policy upon their occurrence. Unlike existing few-shot visual perception tasks [12], failures can result in significant disruptions to surrounding objects and humans in the real world. This nature often restricts real-world experiments to simple tasks. Our AED benchmark is built within Coppeliasim [13] and Pyrep [14], encompassing six indoor tasks and one factory task. This comprehensive benchmark comprises 322 base and 153 new environments, spanning diverse domains and incorporating multiple stages. We aim to create a thorough evaluation platform for AED methodologies, ensuring their effectiveness before real-world deployment.
@@
The AED task present three novel challenges, illustrated in Figure 1. First, AED entails monitoring a policy’s behavior in novel environments, the normal states of which are not observed during training. Second, detecting behavior errors becomes challenging, as there are no noticeable changes to indicate when such errors occur. Specifically, this makes it difficult to discern either minor visual differences or misunderstandings of the task through adjacent frames. Third, AED requires online detection to terminate the policy timely, lacking complete temporal information of the rollout.
@@
Current approaches struggle to address the unique challenges posed by AED. One-class classification (OCC) methods, including one-class SVMs [15] or autoencoders [16, 17, 18, 19], face difficulties in handling unseen environments. These methods are trained solely with normal samples and identify anomalies by evaluating their significant deviation from the learned distribution. However, normal states in novel environments, which include unseen backgrounds and objects, are already considered out-of-distribution for these techniques, resulting in subpar performance. Multiple instance learning [20] or patch-based methods [21, 22] may alleviate the second challenge, particularly minor visual differences, in seen environments. However, the feasibility of applying these methods to AED remains underexplored. Video few-shot anomaly detection (vFSAD) methods [21, 23, 24] are unsuitable for AED due to their dependency on full temporal observation from videos.
@@
To this end, we introduce Pattern Observer (PrObe), a novel algorithm that extracts discriminative features from the monitored policy to identify instances of behavior errors. Specifically, PrObe designs a gating mechanism to extract task-relevant features from the monitored FSI policy to mitigate the impact of a novel environment (first challenge). Then, we design an effective loss function to distill sparse pattern features, making it easier to observe changes in observation (second challenge). Additionally, we design a recurrent generator to generate a pattern flow of current policy. Finally, we determine if there is a behavior error by proposing a novel temporal-aware contrastive objective to compare the pattern flow and demonstrations (third challenge).
@@
Our work makes three significant contributions: (1) We define a vital yet underexplored task called adaptable error detection (AED). (2) We introduce a novel benchmark and present PrObe, facilitating the research community to delve into this important area collectively. (3) We conduct thorough evaluations on our proposed benchmark, demonstrating the effectiveness of PrObe. It surpasses several baseline models and exhibits robustness across different policies. Our research is anticipated to serve as a key foundation for future real-world experiments in the field of FSI research.
@@ 2.1 Few-shot Imitation (FSI)
With the recent progress in meta-learning [25], the community explores the paradigm for learning policies from limited demonstrations during inference [26, 27, 28]. Notably, these works either assume that the agent and expert have the same configuration or explicitly learn a motion mapping between them [27]. Conversely, DC methods [1, 11, 29] develop a policy that behaves conditioned both on the current state and demonstrations. Furthermore, they implicitly learn the mapping between agents and experts, making fine-tuning optional. Thereby, effectively extracting knowledge from demonstrations becomes the most critical matter. In most FSI works, no environment interactions are allowed before inference. Hence, policies are usually trained by behavior cloning (BC) objectives, i.e., learning the likelihood of expert actions by giving expert observations. Recently, DCRL [1] trains a model using reinforcement learning (RL) objects and performs FSI tasks without interacting with novel environments.
@@
Since humans can perform complex long-horizon tasks after watching a few demonstrations, FSI studies continuously pursue solving long-horizon tasks to verify if machines can achieve the same level. A set of research applies policy on a robot arm to perform daily life tasks in the real world [27] or simulation [7]. The task is usually multi-stage and composed of primitives/skills/stages [11, 2], such as a typical pick-and-place task [28] or multiple boxes stacking [7]. Besides, MoMaRT [19] tackles a challenging mobile robot task in a kitchen scene.
@@ 2.2 Few-shot Anomaly Detection (FSAD)
Most existing FSAD studies deal with anomalies in images [22, 30, 31, 32, 33] and a few tackle anomalies in videos [34, 35, 36, 24]. Moreover, problem settings are diverse. Some works presume only normal data are given during training [30, 33, 34, 35], while others train models with normal and a few anomaly data and include unknown anomaly classes during inference [31, 36].
@@
Studies that only use normal training samples usually develop a reconstruction-based model with auxiliary objectives, such as meta-learning [34], optical flow [35], and adversarial training [30, 35]. Besides, patch-based methods [31, 22] reveal the performance superiority on main image FSAD benchmark [37] since the anomaly are tiny defeats. Regarding video few-shot anomaly detection (vFSAD), existing works access a complete video to compute the temporal information for determining if it contains anomalies. In addition, vFSAD benchmarks [38, 39] provide frame-level labels to evaluate the accuracy of locating anomalies in videos.
@@
Although both the vFSAD and our AED task require methods to perform in unseen environments, there are differences: (1) The anomalies in vFSAD involve common objects, while AED methods monitor the policy’s behavior errors. (2) An anomaly in vFSAD results in a notable change in the perception field, such as a cyclist suddenly appearing on the sidewalk [38]. However, no notable change is evident when a behavior error occurs in AED. (3) The whole video is accessible in vFSAD [21, 23, 24], allowing for the leverage of its statistical information. In contrast, AED requires online detection to terminate the policy timely, lacking the complete temporal information of the rollout. These differences make our AED task more challenging and also render vFSAD methods infeasible.
@@ 3 Preliminaries
FSI is a framework worthy of attention, accelerating the development of various robot applications. Following [11], a FSI task is associated with a set of base environments $E^{b}$ and novel environments $E^{n}$. In each novel environment $e^{n}\in E^{n}$, a few demonstrations $\mathcal{D}^{n}$ are given. The objective is to seek a policy $\pi$ that achieves the best performance (e.g., success rate) in novel environments leveraging a few demonstrations. Note that, a task in base and novel environments are semantically similar, but their backgrounds and interactive objects are disjoint. The framework takes as input $N$ demonstrations (collected by a RGB-D camera) and an RGB-D image of the current observation, following the setting of [10, 11]. Addressing FSI tasks typically involves three challenges in practice [11]: (1) The task is long-horizon and multi-stage. (2) The demonstrations are length-variant, making each step misaligned, and (3) The expert and agent have a distinct appearance or configuration. Developing a policy to solve the FSI task and simultaneously tackle these challenges is crucial.
@@
4 Adaptable Error Detection (AED)
Our AED task is formulated to monitor and report FSI policies’ behavior errors, i.e., states of policy rollouts that are inconsistent with demonstrations. The challenges posed by the AED task have been presented in Figure 1. We formally state the task and describe the protocol we utilized below.
Task statement
Let $c$ denote the category of agent’s behavior status, where $c\in\mathbf{C},\mathbf{C}=\{\textrm{normal},\textrm{error}\}$. When the agent with the trained FSI policy $\pi_{\theta}$ performs in a novel environment $e^{n}$, an adaptable error detector $\phi$ can access the agent’s rollout history $h$ and few expert demonstrations $\mathcal{D}^{n}$. It then predicts the categorical probability $\hat{y}$ of the behavior status for the latest state in the history $h$ by $\hat{y}=\phi(h,\mathcal{D}^{n})=P(c\mid enc(h,\mathcal{D}^{n})),$ where $enc$ denotes the feature encoder, and it may be $enc_{\phi}$ ($\phi$ contains its own encoder) or $enc_{\pi_{\theta}}$ (based on policy’s encoder). Next, let $y$ represent the ground truth probability, we evaluate $\phi$ via the expectation of detection accuracy over agent rollouts $X^{n}$ in all novel environments $E^{n}$:
@@ 4 Adaptable Error Detection (AED)
where $A(\cdot,\cdot)$ is the accuracy function that returns 1 if $\hat{y}$ is consistent with $y$ and 0 otherwise. However, frame-level labels are often lacking in novel environments due to the established assumption [8, 11] that we have less control over them. Therefore, we employ a sequence-level $A(\cdot,\cdot)$ in our experiments.
Figure 2: Our AED protocol. the successful agent rollouts $X^{b}_{succ}$, failed agent rollouts $X^{b}_{fail}$, and a few expert demonstrations $\mathcal{D}^{b}$ are available for all base environments $E^{b}$. Then, the task contains three phases: policy training, AED training, and AED inference. We aim to train an adaptable error detector $\phi$ to report policy $\pi_{\theta}$’s behavior errors when performing in novel environments $E^{n}$.
@@
We explain the utilized protocol (shown in Figure 2) with a practical scenario: A company develops a home robot assistant. This robot can perform a set of everyday missions in customers’ houses (novel environments $E^{n}$), given a few demonstrations $\mathcal{D}^{n}$ from the customer. Before selling, the robot is trained in the base environments $E^{b}$ built by the company. In this scenario, both the agent’s and expert’s optimal policies $\pi^{\ast}$ are available during training, with an established assumption from the FSI society [8, 10, 11] that base environments are high controllable. Then, we can collect successful agent rollouts $X^{b}_{succ}$ and a few expert demonstrations $\mathcal{D}^{b}$ for all base environments  [^1] The successful agent rollouts are not expert demonstrations since thet might have different configurations.. Next, we also collect failed agent rollouts $X^{b}_{fail}$ by intervening in the agent’s $\pi^{\ast}$ at a critical timing (e.g., the moment to grasp objects) so that $X^{b}_{fail}$ can possess precise frame-level error labels.
With these resources, our utilized AED protocol consists of three phases: (1) The policy $\pi_{\theta}$ is optimized using successful agent rollouts $X^{b}_{succ}$ and few demonstrations $\mathcal{D}^{b}$. (2) The adaptable error detector $\phi$ is trained on agent rollouts $X^{b}_{succ}$, $X^{b}_{fail}$ and few demonstrations $\mathcal{D}^{b}$. Besides, the detector $\phi$ may use features extracted from policy $\pi_{\theta}$’s encoder, whose parameters are not updated in this phase. (3) The policy $\pi_{\theta}$ solves the task leveraging few demonstrations $\mathcal{D}^{n}$, and the detector $\phi$ monitors the policy’s behavior simultaneously. Notably, no agent rollouts are collected in this phase. Only a few demonstrations $\mathcal{D}^{n}$ are available since the agent is now in novel environments $E^{n}$.